Using Greens theorm ..integration (3x+4y)dx+(2x-3y)dy for a circle




1
Expert's answer
2022-04-07T14:40:23-0400

ANSWER

Let

"C= \left \{ \left ( x,y \right )\in R^{2} :\left ( x-a \right )^{2}+\left ( y-b \right )^{2}=r^{2}\right \}"

be a positively oriented circle and let

"D= \left \{ \left ( x,y \right )\in R^{2} :\left ( x-a \right )^{2}+\left ( y-b \right )^{2}\leq r^{2}\right \}" .

Since the partial derivatives of the functions "L(x,y)=3x+4y," "M(x,y)=2x-3y" are continuous , then by Green's Theorem

"\oint _{C}\left ( Ldx+Mdy \right )=\iint_{D} \left ( \frac{\partial M}{\partial x}-\frac{\partial L}{\partial y } \right )dxdy"

"\oint _{C}\left ( (3x+4y )dx+(2x-3y)dy \right )=\iint_{D} \left ( 2 -4 \right )dxdy=-2(area\: of \: D)=-2\pi r^{2}"

(because "\frac{\partial(2x-3y) }{\partial x}=2, \: \: \frac{\partial(3x+4y) }{\partial y}=4, \: \:" and the area of a circle "D" is equal to "\pi r^{2}" )